find 2 numbers whose sum is 78 and product is 1512

Answer by Alan3354(69443)

78/2 = 39
39^2 = 1521
1521 - 1512 = 9
sqrt(9) = 3
--> 39 - 3 & 39 + 3
--> 36 & 42
--------------
the usual approach:
n(78-n) = 1512
78n - n^2 = 1512
n^2 - 78n + 1512 = 0
Now you have to either factor 1512, or use the quadratic equation.
(n - 36)*(n - 42) = 0
--> 36 & 42